A company got approval to commercialize a new drug to treat Shingles. The following are the key facts: • If the drug passes sPlease answer the question step by step. Thank you!
Answer #1

Solution:-

To calculate NPV, we must calculate the expected annual cash flows if stage 3 trials are a success as follows:

Expected annual cash inflows= (Annual cash inflow if a commercial success*Probability of commercial success) + (Annual cash inflow if not a commercial success*Probability of not being a commercial success)

Expected annual cash inflows= ($6m*40%) + ($1.5m*60%)= $2.4m + $0.9m = $3.3m

Net present value (NPV)= Expected annual cash inflows*Sum of PV factors for 20 years

Sum of PV factors for 20 years (@10%) = 1/1.1 + 1/(1.1)2 + 1/(1.1)3 + 1/(1.1)4 +...…………..+ 1/(1.1)20

Sum of PV factors for 20 years (@10%) = 8.514

Net present value (NPV)= Expected annual cash inflows*Sum of PV factors for 20 years

Net present value (NPV)= $3.3m*8.514 = $28 million